# Obfuscation-Resilient Binary Code Similarity Analysis using Dominance Enhanced Semantic Graph

### v1

@@ 19       Split the target operand of $inst$ as a node $op$;
22             Split $op$ as a node;
23             $\mathcal{G}_{\text{TSG}}$.addDataEdge($opc$, $op$);
@@ 26 $\mathcal{T}_{\textnormal{PD}}\leftarrow\textnormal{getPostDominatorTree}(\mathcal{G}_{\text{CFG}})$;
28       Create a virtual basic block node $bb_{i}$ for $block$;
@@
32       $\mathcal{G}_{\text{BBSG}}$.addDomainceEdge($bb_{j}$, $bb_{i}$);
@@
36      $\mathcal{G}_{\text{BBSG}}$.addPostDominanceEdge($bb_{j}$, $bb_{i}$);
@@ 3.3. Model Training
GGNN (Li et al., 2015) is applied to generate embedding. We first present the GGNN’s structure before discussing the model training.
@@ 3.3.1. GGNN
GNN is a neural network model for encoding nodes and edges, and learning their embeddings. GNN consists of multiple message passing layers and a graph pooling layer. GGNN is a variant of GNN, which extends GNN by incorporating a Gated Recurrent Unit (GRU) activation in the message passing phase to capture information from neighboring nodes. Following (He et al., 2024; Li et al., 2023; Guo et al., 2022), we adopt GGNN to learn the DESG’s semantics.
@@
(1) $\mathbf{m}_{u}^{l}=\sum_{v\in\text{Edge}(u)}f\left(\mathbf{h}_{u}^{l},\mathbf{h}_{v}^{l},\mathbf{e}_{v,u},\mathbf{e}_{u,v}\right)$
@@
(2) $\mathbf{h}_{u}^{l+1}=\text{GRU}\left(\mathbf{h}_{u}^{l},\mathbf{m}_{u}^{l}\right)$
@@
Graph Pooling. After message passing layers, the final step is to pool all node embeddings into a graph embedding. Softmax (Li et al., 2020) is a widely used pooling method, which is formulated as Eqn. (3).
(3) $\text{Softmax}(\mathcal{H}\mid\beta)=\sum_{\mathbf{h}_{u}^{L}\in\mathcal{H}}\frac{\exp\left(\beta\odot\mathbf{h}_{u}^{L}\right)}{\sum_{\mathbf{h}_{v}^{L}\in\mathcal{H}}\exp\left(\beta\odot\mathbf{h}_{v}^{L}\right)}\odot\mathbf{h}_{u}^{L}$
where $\mathcal{H}$ refers to the set of all node hidden states and $\beta$ is a parameter known as the inverse temperature. Multi-head attention is a core technique in the transformer architecture (Vaswani, 2017), allowing each head to learn different attention distributions. Similar to (He et al., 2024), we integrate softmax pooling with multi-head attention, forming a stack of multiple softmax layers. We first transform the hidden state $\mathbf{x}_{u}^{k}=\text{LayerNorm}\big{(}\text{ReLU}(\text{Linear}_{k}(\mathbf{h}_{u}^{r}))\big{)}$ using $k$ linear layers, followed by a Relu activation and layer normalization. Then, the calculation of every head is formulated as Eqn. (4).
(4) $\mathbf{e}_{g}^{k}=\text{Softmax}\left(\left\{\mathbf{x}_{u}^{k}\mid u\in\mathcal{N}\right\}\mid\beta^{k}\right)$
@@ 3.3.2. Model Training
We employ distance weighted negative sampling (Wu et al., 2017) with a margin-based pairwise loss (Li et al., 2019) to train the GGNN. This approach helps the model effectively distinguish between positive and negative samples, enhancing the capability of the embeddings. The loss function is formulated as Eqn. (5).
(5) $\min_{\theta}\mathcal{L}(\theta)=\sum_{\langle\mathbf{p},\mathbf{q}\rangle\in\mathcal{B}}\big{(}\text{positive}(\mathbf{p},\mathbf{q})+\text{negative}(\mathbf{p})\big{)}$
@@ 4.1. Dataset
Our dataset is built on the BinKit dataset (Kim et al., 2023), a large-scale benchmark for BCSA. It includes normal and obfuscation versions generated using the same optimization levels and ISAs. Normal versions include 3 ISAs (MIPS, ARM, and x86), 9 compiler versions (5 GCC, 4 Clang versions), and 5 different optimization levels (O0 to O3 and Os). Based on Obfuscator-LLVM (O-LLVM) (Junod et al., 2015), obfuscation versions includes 4 obfuscation options: instruction substitution (SUB), bogus control flow (BCF), control flow flattening (FLA), and a combination of all obfuscation options (SUB+BCF+FLA), referred to as ALL. To evaluate the effectiveness of ORCAS in analyzing the semantic similarity of binary functions across different obfuscation options, optimization levels, and ISAs, we integrate binaries from both normal and obfuscation versions. For our experiments, we select binaries from the x86-64 and ARM-64 architectures, compiled with Clang for normal versions and O-LLVM (BCF, FLA, SUB, and ALL) for obfuscation versions, specifically with O0 and O3 optimization levels, covering 578 binaries.
@@ 4.2.1. One-to-One Similarity Matching
Several works (Haq and Caballero, 2021) utilize accuracy or recall to evaluate the performance of the model in similarity matching. However, these metrics at a fixed threshold may not fully reflect a model’s ability to distinguish similar and dissimilar pairs. Instead, we evaluate and compare the similarity matching performance of different approaches using the Precision-Recall (PR) curve. Let $TP$, $FP$, and $FN$ represent the number of True Positive, False Positive, and False Negative binary function pairs inferred by the model at a given threshold, respectively, we have $\text{Precision}={TP/(TP+FP)}$ and $\text{Recall}={TP/(TP+FN)}$. To save space, we employ the the Area Under the PR Curve (PR-AUC) as a comprehensive metric to summarize the overall performance, providing a more comprehensive comparison across different approaches.
@@ 4.2.2. One-to-Many Similarity Searching
(6) $\text{Recall}@k={1\over n}\sum_{i=1}^{n}{\mathbb{I}\big{(}{\text{Rank}(f_{i}^{gt})}\leq k\big{)}}$
@@
(7) $\text{MRR}={1\over n}\sum_{i=1}^{n}{1\over{\text{Rank}(f_{i}^{gt})}}$
@@ 4.3. Baselines
SAFE (Massarelli et al., 2022) employs a recurrent neural network (RNN) augmented with attention mechanisms. It generates the embeddings of binary functions from the assembly instructions.
Gemini (Xu et al., 2017b) takes the CFG as input, assigning manually crafted features to each basic block to construct an attributed CFG. It then uses GNN to calculate the similarity between binary functions by introducing a Siamese network.
Asm2vec (Ding et al., 2019) uses random walks model the CFG as multiple sequences and applies the PV-DM model (Le and Mikolov, 2014) to learn the embeddings of binary function.
Graph Matching Networks (GMN) (Li et al., 2019) proposes a variant of the GNN that incorporates a cross-graph attention-based matching mechanism. It takes pairs of CFG as input and computes a similarity score between them.
Trex (Pei et al., 2023) first employs a hierarchical transformer-based model to capture the execution semantics of binary functions from micro traces. Then, it transfers the acquired semantic knowledge to match semantically similar binary functions. Our implementation is directly based on the source code released by the authors. [^3] https://github.com/CUMLSec/trex
HermesSim (He et al., 2024) introduces novel SOG to represent binary function because it has well-defined semantic structures. Then, it utilizes GGNN to generate the embeddings of binary functions and calculate their cosine similarity. Our implementation is based on the source code released by the authors. [^4] https://github.com/NSSL-SJTU/HermesSim
jTrans (Wang et al., 2022) proposes a jump-aware transformer-based model pre-trained on a large dataset, BinaryCorp, comprising approximately 21 million x86 binary functions. We fine-tune the pre-trained model released by the anthors on our dataset. [^5] https://github.com/vul337/jTrans
CRABS-former (Feng et al., 2024) is a BCSA model built on top of jTrans. It introduces several assembly normalization strategies for x86-64 and ARM-64 to improve the performance of BCSA. We implement CRABS-former and fine-tune it on our dataset.
Our implementation of SAFE, Gemini, Asm2vec, and GMN is based on the implementation provided by (Marcelli et al., 2022), with the default parameter settings. [^6] https://github.com/Cisco-Talos/binary_function_similarity
@@ 5.2. RQ2: Matching Obfuscated Binary Functions
Similar to (Ding et al., 2019; Zhang et al., [n. d.]; Pei et al., 2023; Feng et al., 2024), we evaluate ORCAS’s ability to match obfuscated binary functions with their normal version. First, we select representative projects (i.e., Readline, Sed, Sharutils, and Tar) from our dataset. For each project, we compile all functions without applying any obfuscation techniques. Next, we compile the project again using the selected obfuscation technique. Then, we can obtain both normal and obfuscated versions for each function. We link the normal and obfuscated functions to establish a one-to-one match using function symbols for evaluation. We consider function pairs compiled from the same function as positive samples, with others treated as negative samples, setting a 1:100 ratio of positive to negative samples. Finally, we use the PR-AUC score as the evaluation metric for each project.
@@ 5.4. RQ4: Real-World Vulnerability Search
Previous works (Luo et al., 2023; Wang et al., 2022, 2024) utilize 1-day vulnerability detection to evaluate the effectiveness of BCSA approaches in real-world tasks. Similar to (Wang et al., 2022), we perform our evaluation on eight Common Vulnerabilities and Exposures (CVEs) from a public dataset presented in (Wang et al., 2019). We produce 10 variants for each function by O-LLVM (Junod et al., 2015) with 2 ISAs (x86-64 and ARM-64) and 5 obfuscation options (disable obfuscation, enable SUB, BCF, FLA, and ALL obfuscation options). We enable -O3 optimization option to make them closer to the release version as suggested in (Kim et al., 2023), and apply -fno-inline-functions option to facilitate locating a certain vulnerable function. Thus, we obtain a challenging obfuscated real-world vulnerability dataset, which has been released to facilitate future research on obfuscated binary code analysis. [^7] https://anonymous.4open.science/r/ORCAS-CF05
We provide details for evaluating the effectiveness of vulnerability detection. Similar to (Wang et al., 2022, 2024), for each CVE, we first extract the functions in the 10 project variants to build a function pool $P$ and mark the 10 vulnerable functions in $P$. Then, we search in $P$ using a variant of the vulnerable function and it returns the top 10 similar binary functions. The number of function pools for each CVE varies from 18,719 to 57,673. Finally, assuming $m$ of the returned binary functions are marked, we calculate the recall rate by $m/10$.
@@ 6.1. Binary Code Representation in BCSA
Trees. TEDEM (Pewny et al., 2014) transforms basic block instructions into equations through symbolic simplification to generate expression trees, then measures similarity using tree edit distance. Recent approach by Asteria (Yang et al., 2021) finds that abstract syntax tree (AST) exhibit greater stability than CFG cross-ISA, employing Tree-LSTM networks to encode AST of a binary function into embeddings. Building on Asteria, Asteria-Pro (Yang et al., 2023a) enhances detection efficiency by incorporating pre-filtering and re-ranking modules. RCFG2Vec (Li et al., 2024) converts each assembly instruction into a three-level syntax tree, where the root represents the instruction itself and child nodes denote its opcode or operands, subsequently generating three types of embeddings.
Instruction streams. Both assembly code and IR code are widely adopted for binary representation (Massarelli et al., 2022; Li et al., 2021; Wang et al., 2022; Ahn et al., 2022; Luo et al., 2023). Early approach SAFE (Massarelli et al., 2022) employed skip-gram models, which predict surrounding instructions from a target instruction, to generate instruction embeddings. With advances in NLP technologies, pre-trained models have demonstrated superior performance. Researchers (Li et al., 2021; Wang et al., 2022; Ahn et al., 2022; Luo et al., 2023) develop specialized pre-training tasks to enhance semantic capture, with PalmTree (Li et al., 2021) implementing both context window prediction and def-use prediction tasks. Moreover, Vulhark (Luo et al., 2023) further enhancing IR through microcode-based optimizations including root-operand analysis and instruction simplification. However, substituting complex instructions for simpler ones significantly degrades representation effectiveness under obfuscation.
Graphs. CFG serve as a prevalent representation owing to their cross-ISA ability (He et al., 2024), often extended through integration with other features. Early approaches (Feng et al., 2016; Xu et al., 2017b) construct a attributed CFG using manually designed basic block features, while recent approaches integrate CFG with CG and DFG (Kim et al., 2022; Duan et al., 2020; Guo et al., 2022). DeepBinDiff (Duan et al., 2020) merges CG and CFG to construct a inter-procedural CFG. XBA (Kim et al., 2022) propose a binary disassembly graph that leverages basic blocks and external functions. Additionally, HermesSim (He et al., 2024) proposes a SOG that decomposes instructions into operand and opcode nodes to reveal their relations. These approaches achieve promising performance but become unstable under obfuscation due to their reliance on original control flow.

### v2

@@ 19  Split the target operand of instinst as a node opop;
22     Split opop as a node;
23     $\mathcal{G}_{\text{TSG}}$.addDataEdge(opcopc, opop);
@@ 26 $\mathcal{T}_{\textnormal{PD}}\leftarrow\textnormal{getPostDominatorTree}(\mathcal{G}_{\text{CFG}})$;
28  Create a virtual basic block node $bb_{i}$ for blockblock;
@@
32  $\mathcal{G}_{\text{BBSG}}$.addDomainceEdge($b_{j}, bb$$b_{i}$);
@@
36 $\mathcal{G}_{\text{BBSG}}$.addPostDominanceEdge($b_{j}, bb$$b_{i}$);
@@ 3.3. Model Training
GGNN (GGNN2015) is applied to generate embedding. We first present the GGNN’s structure before discussing the model training.
@@ 3.3.1. GGNN
GNN is a neural network model for encoding nodes and edges, and learning their embeddings. GNN consists of multiple message passing layers and a graph pooling layer. GGNN is a variant of GNN, which extends GNN by incorporating a Gated Recurrent Unit (GRU) activation in the message passing phase to capture information from neighboring nodes. Following (sog2024; sense2023; guo2022), we adopt GGNN to learn the DESG’s semantics.
@@
(1) $\footnotesize\mathbf{m}_{u}^{l}=\sum_{v\in\text{Edge}(u)}f\left(\mathbf{h}_{u}^{l},\mathbf{h}_{v}^{l},\mathbf{e}_{v,u},\mathbf{e}_{u,v}\right)$
@@
(2) $\footnotesize\mathbf{h}_{u}^{l+1}=\text{GRU}\left(\mathbf{h}_{u}^{l},\mathbf{m}_{u}^{l}\right)$
@@
Graph Pooling. After message passing layers, the final step is to pool all node embeddings into a graph embedding. Softmax (softmax2020) is a widely used pooling method, which is formulated as Eqn. (3).
(3) $\footnotesize\text{Softmax}(\mathcal{H}\mid\beta)=\sum_{\mathbf{h}_{u}^{L}\in\mathcal{H}}\frac{\exp\left(\beta\odot\mathbf{h}_{u}^{L}\right)}{\sum_{\mathbf{h}_{v}^{L}\in\mathcal{H}}\exp\left(\beta\odot\mathbf{h}_{v}^{L}\right)}\odot\mathbf{h}_{u}^{L}$
where $\mathcal{H}$ refers to the set of all node hidden states and $\beta$ is a parameter known as the inverse temperature. Multi-head attention is a core technique in the transformer architecture (vaswani2017attention), allowing each head to learn different attention distributions. Similar to (sog2024), we integrate softmax pooling with multi-head attention, forming a stack of multiple softmax layers. We first transform the hidden state $\mathbf{x}_{u}^{k}=\text{LayerNorm}\big{(}\text{ReLU}(\text{Linear}_{k}(\mathbf{h}_{u}^{r}))\big{)}$ using $k$ linear layers, followed by a Relu activation and layer normalization. Then, the calculation of every head is formulated as Eqn. (4).
(4) $\footnotesize\mathbf{e}_{g}^{k}=\text{Softmax}\left(\left\{\mathbf{x}_{u}^{k}\mid u\in\mathcal{N}\right\}\mid\beta^{k}\right)$
@@ 3.3.2. Model Training
We employ distance weighted negative sampling (wu2017sampling) with a margin-based pairwise loss (GMN2019) to train the GGNN. This approach helps the model effectively distinguish between positive and negative samples, enhancing the capability of the embeddings. The loss function is formulated as Eqn. (5).
(5) $\footnotesize\min_{\theta}\mathcal{L}(\theta)=\sum_{\langle\mathbf{p},\mathbf{q}\rangle\in\mathcal{B}}\big{(}\text{positive}(\mathbf{p},\mathbf{q})+\text{negative}(\mathbf{p})\big{)}$
@@ 4.1. Dataset
Our dataset is built on the BinKit dataset (kim2022revisiting), a large-scale benchmark for BCSA. It includes normal and obfuscation versions generated using the same optimization levels and ISAs. Normal versions include 3 ISAs (MIPS, ARM, and x86), 9 compiler versions (5 GCC, 4 Clang versions), and 5 different optimization levels (O0 to O3 and Os). Based on Obfuscator-LLVM (O-LLVM) (obfuscator2015), obfuscation versions includes 4 obfuscation options: instruction substitution (SUB), bogus control flow (BCF), control flow flattening (FLA), and a combination of all obfuscation options (SUB+BCF+FLA), referred to as ALL. To evaluate the effectiveness of ORCAS in analyzing the semantic similarity of binary functions across different obfuscation options, optimization levels, and ISAs, we integrate binaries from both normal and obfuscation versions. For our experiments, we select binaries from the x86-64 and ARM64 architectures, compiled with Clang for normal versions and O-LLVM (BCF, FLA, SUB, and ALL) for obfuscation versions, specifically with O0 and O3 optimization levels, covering 578 binaries.
@@ 4.2.1. One-to-One Similarity Matching
Several works (ASurveyofBinaryCodeSimilarity2021) utilize accuracy or recall to evaluate the performance of the model in similarity matching. However, these metrics at a fixed threshold may not fully reflect a model’s ability to distinguish similar and dissimilar pairs. Instead, we evaluate and compare the similarity matching performance of different approaches using the Precision-Recall (PR) curve. Let $TP$, FPFP, and $FN$ represent the number of True Positive, False Positive, and False Negative binary function pairs inferred by the model at a given threshold, respectively, we have $\text{Precision}={TP/(TP+FP)}$ and $\text{Recall}={TP/(TP+FN)}$. To save space, we employ the the Area Under the PR Curve (PR-AUC) as a comprehensive metric to summarize the overall performance, providing a more comprehensive comparison across different approaches.
@@ 4.2.2. One-to-Many Similarity Searching
(6) $\footnotesize\text{Recall}@k={1\over n}\sum_{i=1}^{n}{\mathbb{I}\big{(}{\text{Rank}(f_{i}^{gt})}\leq k\big{)}}$
@@
(7) $\footnotesize\text{MRR}={1\over n}\sum_{i=1}^{n}{1\over{\text{Rank}(f_{i}^{gt})}}$
@@ 4.3. Baselines
SAFE (SAFE2022) employs a recurrent neural network (RNN) augmented with attention mechanisms. It generates the embeddings of binary functions from the assembly instructions.
Gemini (gemini2017) takes the CFG as input, assigning manually crafted features to each basic block to construct an attributed CFG. It then uses GNN to calculate the similarity between binary functions by introducing a Siamese network.
Asm2vec (asm2vec2019) uses random walks model the CFG as multiple sequences and applies the PV-DM model (le2014distributed) to learn the embeddings of binary function.
Graph Matching Networks (GMN) (GMN2019) proposes a variant of the GNN that incorporates a cross-graph attention-based matching mechanism. It takes pairs of CFG as input and computes a similarity score between them.
Trex (Trex2023) first employs a hierarchical transformer-based model to capture the execution semantics of binary functions from micro traces. Then, it transfers the acquired semantic knowledge to match semantically similar binary functions. Our implementation is directly based on the source code released by the authors. [^3] https://github.com/CUMLSec/trex
jTrans (jtrans2022) proposes a jump-aware transformer-based model pre-trained on a large dataset, BinaryCorp, comprising approximately 21 million x86 binary functions. We fine-tune the pre-trained model released by the anthors on our dataset. [^4] https://github.com/vul337/jTrans
CRABS-former (crabs-former2024) is a BCSA model built on top of jTrans. It introduces several assembly normalization strategies for x86-64 and ARM64 to improve the performance of BCSA. We implement CRABS-former and fine-tune it on our dataset.
HermesSim (sog2024) introduces novel SOG to represent binary function because it has well-defined semantic structures. Then, it utilizes GGNN to generate the embeddings of binary functions and calculate their cosine similarity. Our implementation is based on the source code released by the authors. [^5] https://github.com/NSSL-SJTU/HermesSim
Our implementation of SAFE, Gemini, Asm2vec, and GMN is based on the implementation provided by (marcelli2022machine), with the default parameter settings. [^6] https://github.com/Cisco-Talos/binary_function_similarity
@@ 5.2. RQ2: Matching Obfuscated Binary Functions
Similar to (asm2vec2019; gtrans2024; Trex2023; crabs-former2024), we evaluate ORCAS’s ability to match obfuscated binary functions with their normal version. First, we select representative projects (i.e., Readline, Sed, Sharutils, and Tar) from our dataset. For each project, we compile all functions without applying any obfuscation techniques. Next, we compile the project again using the selected obfuscation technique. Then, we can obtain both normal and obfuscated versions for each function. We link the normal and obfuscated functions to establish a one-to-one match using function symbols for evaluation. We consider function pairs compiled from the same function as positive samples, with others treated as negative samples, setting a 1:100 ratio of positive to negative samples. Finally, we use the PR-AUC score as the evaluation metric for each project.
@@ 5.4. RQ4: Real-World Vulnerability Search
Previous works (vulhawk2023; jtrans2022; cebin2024) utilize 1-day vulnerability detection to evaluate the effectiveness of BCSA approaches in real-world tasks. Similar to (jtrans2022), we perform our evaluation on eight Common Vulnerabilities and Exposures (CVEs) from a public dataset presented in (wang2019detecting). We produce 10 variants for each function by O-LLVM (obfuscator2015) with 2 ISAs (x86-64 and ARM64) and 5 obfuscation options (disable obfuscation, enable SUB, BCF, FLA, and ALL obfuscation options). We enable -O3 optimization option to make them closer to the release version as suggested in (kim2022revisiting), and apply -fno-inline-functions option to facilitate locating a certain vulnerable function. Thus, we obtain a challenging obfuscated real-world vulnerability dataset, which has been released to facilitate future research on obfuscated binary code analysis. [^7] https://github.com/Cao-Wuhui/ORCAS
We provide details for evaluating the effectiveness of vulnerability detection. Similar to (jtrans2022; cebin2024), for each CVE, we first extract the functions in the 10 project variants to build a function pool $P$ and mark the 10 vulnerable functions in $P$. Then, we search in $P$ using a variant of the vulnerable function and it returns the top 10 similar binary functions. The number of function pools for each CVE varies from 18,719 to 57,673. Finally, assuming $m$ of the returned binary functions are marked, we calculate the recall rate by $m/10$.
@@ 6.1. Binary Code Representation in BCSA
Trees. TEDEM (pewny2014leveraging) transforms basic block instructions into equations through symbolic simplification to generate expression trees. Recent approach by Asteria (asteria2021; asteria2024) finds that abstract syntax tree (AST) exhibit greater stability than CFG cross-ISA, employing Tree-LSTM networks to encode AST of a binary function into embeddings. RCFG2Vec (rcfg2vec2024) converts each assembly instruction into a three-level syntax tree, where the root represents the instruction itself and child nodes denote its opcode or operands, subsequently generating three types of embeddings.
Instruction streams. Both assembly code and IR code are widely adopted for binary representation (SAFE2022; palmtree2021; jtrans2022; binshot2022; vulhawk2023). Early approach SAFE (SAFE2022) employed skip-gram models, which predict surrounding instructions from a target instruction, to generate instruction embeddings. With advances in NLP technologies, pre-trained models have demonstrated superior performance. Researchers (palmtree2021; jtrans2022; binshot2022; vulhawk2023) develop specialized pre-training tasks to enhance semantic capture, with PalmTree (palmtree2021) implementing both context window prediction and def-use prediction tasks. Moreover, Vulhark (vulhawk2023) further enhancing IR through microcode-based optimizations including root-operand analysis and instruction simplification. However, substituting complex instructions for simpler ones significantly degrades representation effectiveness under obfuscation.
Graphs. CFG serve as a prevalent representation owing to their cross-ISA ability (sog2024), often extended through integration with other features. Early approaches (genius2016; gemini2017) construct a attributed CFG using manually designed basic block features, while recent approaches integrate CFG with CG and DFG (kim2022; deepbindiff2020; guo2022). DeepBinDiff (deepbindiff2020) merges CG and CFG to construct a inter-procedural CFG. XBA (kim2022) propose a binary disassembly graph that leverages basic blocks and external functions. Additionally, HermesSim (sog2024) proposes a SOG that decomposes instructions into operand and opcode nodes to reveal their relations. These approaches achieve promising performance but become unstable under obfuscation due to their reliance on original control flow.